OpenStudy (anonymous):

A theater has 600 tickets to sell for a show. Two hundred twenty-five adult tickets sell at $2 more than the student tickets. All the tickets are sold, for a total of $2,250. What is the price of each type of ticket? A. Adult tickets sell at $7, and student tickets sell at $5. B. Adult tickets sell at $6, and student tickets sell at $4. C. Adult tickets sell at $5, and student tickets sell at $3. D. Adult tickets sell at $5, and student tickets sell at $7.

OpenStudy (anonymous):

Can anyone please help?

OpenStudy (whpalmer4):

Let # of adult tickets be A Let # of student tickets be S Adult tickets sell for p Student tickets sell for p-2 (adult tickets are $2 more than student tickets) 600 tickets total, so A + S = 600 A*p + S*(p-2) = 2250 Solve that by substitution, I think...

OpenStudy (whpalmer4):

You can eliminate at least one answer possibility because the relationship between the price of student and adult tickets is incorrect...

OpenStudy (whpalmer4):

You can also rule out 2 choices by asking how much money would come in if the cheapest tickets were sold for all 600 seats. If that amount > $2250, that obviously can't be the answer.

OpenStudy (whpalmer4):

(but you really should solve the equations!)

OpenStudy (anonymous):

ok lets say that x is adults and y is students you do this x + y = 600 and 225x + 2y =2250 ok we are going to use substitution method then do y = -x + 600 plug it into the y in the equation 225x + 2y = 2250 it will be 225x + 2(-x +600) = 2250 then 225x + -2x +1200 = 2250 subtract 1200 from both sides 225x + -2x + 1200 -1200 = 2250 -1200 it will be 225x + -2x = 1050 then add the both x's 223x = 1050 so the x will be 4.7 so round it to $5

OpenStudy (anonymous):

so we found the x amount now we will find the y amount

OpenStudy (whpalmer4):

no rounding necessary, the problem solves exactly!

OpenStudy (anonymous):

Yes I am solving it now, I get it

OpenStudy (anonymous):

the y amount is 3

OpenStudy (anonymous):

Am I correct

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thanks a lot!!!

OpenStudy (anonymous):

so it is c

OpenStudy (whpalmer4):

225 adult tickets sold at price p \[225*p + (600-225)*(p-2) = 2250\]\[225p + 375p-750=2250\]\[600p = 3000\]\[p = 5\] price of adult ticket is 5 price of student ticket is 5-2 = 3

OpenStudy (whpalmer4):

@FriedRice you came up with a funky equation here: 225x + 2y = 2250

OpenStudy (anonymous):

no because you are substituting for y so you need to put it and then add -x + 600

OpenStudy (anonymous):

but you are right because you are a mentor

OpenStudy (whpalmer4):

what do x and y represent in that equation?

OpenStudy (anonymous):

x represents adults and y represents students

OpenStudy (anonymous):

it says that in my equation

OpenStudy (whpalmer4):

okay, what do the coefficients represent? 225 and 2

OpenStudy (anonymous):

Two hundred twenty-five adult tickets sell at $2 more than the student tickets

OpenStudy (whpalmer4):

ah, but that's not what you have. x is the number of adults, so you can't have 225 in front x! it should be price * count

OpenStudy (whpalmer4):

adult tickets sell for the price of student tickets + $2. so, x*(student price + 2) + y*(student price) = 2250

OpenStudy (anonymous):

ohh ok i might have brain farted on that one

OpenStudy (whpalmer4):

we know that x+y=600, so we can substitute (600-x) for y, to give us x(student price +2) + (600-x)(student price) = 2250 but we already know that x = 225 and y = 600-225, so that becomes 225(student price + 2) + 375(student price) = 2250 225 * sp + 225*2 + 375*sp = 2250 600*sp = 2250-450 600*sp = 1800 sp = 3

OpenStudy (whpalmer4):

adult price = student price + 2 = 3 + 2 = 5

OpenStudy (anonymous):

ok thanks for teaching me

OpenStudy (whpalmer4):

it helps sometimes to do problems with the units attached: 3 adults * $5/adult + 2 students * $3/student etc. and make sure that the units come out properly. that would have exposed the faulty equation immediately because you would have had square adults and so on, whereas everyone knows that they are round :-)

OpenStudy (anonymous):

I wish I was as smart as you guys

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

you will become knowledgeable while you grow up and learn more